OpenStudy (anonymous):

how to solve this. find the complementary solution, Yc to the homogeneous ODE d^4y/dx^4 - 4 d^3y/dx^3 + 7 d^2y/dy^2 - 6 dy/dx + 2y = 0

OpenStudy (amistre64):

first, write it in a notation that is flattering

OpenStudy (amistre64):

y^4' - y'''+ 7y'' - 6y' + 2y = 0

OpenStudy (amistre64):

then root it out like any other equation

OpenStudy (amistre64):

\[r^4 - r^3 +7r^2-6r+2=0\] such that:\[y=c_1e^{r_1x}+c_2e^{r_2x}+c_3e^{r_3x}+c_4e^{r_4x}\]

OpenStudy (amistre64):

start making some educated guesses for a root :) 1,-1,2,-2 seem to be the best options for a rational solution

OpenStudy (anonymous):

how you get that root

OpenStudy (amistre64):

what root? i havent found a root yet, just some educated guesses at what it might be to try out before things get hairy

OpenStudy (amistre64):

1 -1 7 -6 2 0 -2 6 1 -3 13 ...... the rational roots doesnt seem to wanna play nice

OpenStudy (amistre64):

im missing a 4 in the equation .... thanks to the hideous way it was entered

OpenStudy (amistre64):

\[r^4 - 4r^3 +7r^2-6r+2=0\]

OpenStudy (amistre64):

now 1 IS a root :) when r=1 1 - 4 +7- 6+2 = 10-10 = 0

OpenStudy (amistre64):

1 -4 +7 -6 +2 0 1 -3 4 1 -3 4 -2 .................. \[(r-1)(r^3-3r^2+4r-2)=0\]

OpenStudy (amistre64):

r=1 again; 1-3+4-2 = 5-5 = 0

OpenStudy (amistre64):

1 -3 +4 -2 0 1 -2 1 -2 2 -------- \[(r-1)^2(r^2-2r+2)=0\]

OpenStudy (amistre64):

the rest are complex r^2-2r = -2 (r-1)^2 = -1 r = 1+- i sooo, 2 repeats and a complex

OpenStudy (amistre64):

\[y=c_1e^x+c_2x\ e^x+e^x(c_3cos(x)+c_4sin(x))\]

OpenStudy (amistre64):

now i can answer any questions since we know what the end result should be :)

OpenStudy (anonymous):

is that the final answer

OpenStudy (amistre64):

dunno, feel free to dbl chk my math to make sure ...

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

yw